# Exploiting Kubernetes’ Image Pull Implementation to Deny Node Availability

### v1

@@ VIII Conclusions
This paper demonstrates that the asynchronous implementation of some routes from the CRI-API can be exploited to create a resource-based DoS. We illustrate this threat through three distinct attacks. First, we focus on resource usage, where the attack used up to 95% of the CPU. Second, we show how this attack could interfere with other applications running in the cluster. Finally, we investigate how the cache could be manipulated by the constant stream of requests performed by the attack. We further experimentally prove the effectiveness of each attack in a local testbed and on GKE. Finally, we provide a guideline to solve the problem in all the related projects and are currently collaborating with the working groups to solve this issue as soon as possible. As a stopgap, we present an eBPF-based proof-of-concept to monitor the activity in the cluster, stopping useless image downloads.
@@ Acknowledgments
This work was partially supported by project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU and by European Union’s Horizon Europe Programme under grant agreement No 101070473 (project FLUIDOS).

### v2

@@ VIII Conclusions
This paper demonstrates that the asynchronous implementation of some routes from the CRI-API can be exploited to create a resource-based DoS. We illustrate this threat through three distinct attacks. First, we focus on resource usage, where the attack used up to 95% of the CPU. Second, we show how this attack could interfere with the performance of other applications that run on the affected worker nodes of the cluster. Finally, we investigate how the cache could be manipulated by the constant stream of requests performed by the attack. We further experimentally prove the effectiveness of each attack in a local testbed and on GKE. Finally, we provide a guideline to solve the problem in all the related projects and are currently collaborating with the working groups to solve this issue as soon as possible. As a stopgap, we present an eBPF-based proof-of-concept to monitor the activity in the cluster, stopping useless image downloads.
@@ Acknowledgments
This work was partially supported by project SERICS (PE00000014), MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU, and by project FLUIDOS (grant agreement No 101070473), European Union’s Horizon Europe Programme.